# Movement-Prediction-Adjusted Naïve Forecast

### v8

@@ 2.2 Parameter Optimization Using In-Sample Data
Accordingly, the MSE difference between two types of naïve forecasts is as follows:
$\Delta MSE_{\text{in-sample}}$ $=\text{MSE}_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (7)
@@
where $\theta_{\text{in-sample}}$ represents the parameter $\theta$ in the adjusted naïve forecast for the in-sample set and $\bar{\epsilon}_{\text{in-sample}}$ denotes the mean of $|\epsilon_{t}|$ calculated over the in-sample set. A detailed algebraic derivation of the $\Delta MSE_{\text{in-sample}}$ is provided in Appendix A.
According to Eq. (7), the optimization reduces to maximizing the following function concerning parameter $\theta_{\text{in-sample}}$:
@@
Differentiating Eq. (8) and setting the derivative to zero yields the optimal solution:
Proposition 1 (Optimal parameter solution). The optimal parameter value $\theta_{\text{in-sample}}^{*}$ that maximizes the MSE difference between the adjusted naïve forecast and the original naïve forecast on the in-sample set is given by:
@@
This result explicitly reveals that the optimal value of the parameter $\theta$ depends linearly on the directional accuracy of the movement predictions for a sequence with a fixed length. Since the second derivative of $f(\theta_{\text{in-sample}})$ is negative, this optimal solution is a maximum.
@@
Proposition 2 (Performance guarantee). Under the optimal parameter setting using $\theta_{\text{in-sample}}^{*}$, the adjusted naïve forecast achieves a nonnegative improvement over the original naïve forecast in terms of the MSE on the in-sample set:
@@
Proposition 2 confirms that as the directional accuracy increases beyond 0.5, the value of $\theta_{\text{in-sample}}^{*}$ increases from 0, leading to progressively greater improvements over the original naïve forecast. At a directional accuracy of 1, the maximum achievable improvement is $\bar{\epsilon}_{\text{in-sample}}^{2}$. Thus, with the optimal parameter setting, the adjusted naïve forecast matches or surpasses the performance of the original naïve forecast within the in-sample dataset:
$MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}\leq MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}$ (11)
This result underlines the theoretical robustness of the proposed adjustment methodology.
2.3 Uncertainty Analysis and Practical Generation of Out-of-Sample Forecasts
We now examine the performance of the adjusted naïve forecast via out-of-sample data. Suppose that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are used as the parameters of the adjusted naïve forecast on the out-of-sample set, which contains $M$ time steps. Suppose that the movement predictions on this set yield $m_{\text{out-of-sample}}^{\text{correct}}$ correct and $m_{\text{out-of-sample}}^{\text{incorrect}}$ incorrect predictions; then, the directional accuracy of the movement prediction on the out-of-sample set can be expressed as:
@@ 2.3 Uncertainty Analysis and Practical Generation of Out-of-Sample Forecasts
Consequently, the MSE difference between two types of naïve forecasts on the out-of-sample is expressed as:
$\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (13)
@@
where $\bar{\epsilon}_{\text{out-of-sample}}$ denotes the actual mean absolute increment on the out-of-sample set and $\theta_{\text{out-of-sample}}^{*}$ denotes the actual optimal value of $\theta$ for the out-of-sample set, which is derived according to Proposition 1:
@@
A detailed algebraic expansions of the $\Delta MSE_{\text{out-of-sample}}$ is provided in Appendix B. For the adjusted naïve forecast to outperform the original naïve forecast, the $\Delta MSE_{\text{out-of-sample}}$ should be no less than zero. Therefore, according to Eq. (13), the following inequality must hold:
@@
Given that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are specific estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, re2spectively, inequality (15) can be generalized as:
Proposition 3 (Out-of-sample performance criterion). The necessary condition for the adjusted naïve forecast to achieve an improvement over the original naïve forecast in terms of the MSE on the out-of-sample set is as follows:
@@
where $\hat{\theta}_{\text{out-of-sample}}$ represents the estimate of $\theta_{\text{out-of-sample}}^{*}$ and $\hat{\bar{\epsilon}}_{\text{out-of-sample}}$ represents the estimate of $\bar{\epsilon}_{\text{out-of-sample}}$.
Proposition 3 establishes a clear boundary criterion when an adjusted naïve forecast can outperform the original naïve forecast, specifically depending on whether the product of the actual optimal parameter values exceeds half the product of the estimated optimal parameter values. We can treat $\theta_{\text{out-of-sample}}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ as random variables, each following a distribution defined over nonnegative values. Accordingly, inequality (16) represents a probabilistic event, the likelihood of which depends on the realized values of these variables. Since $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ cannot be known in advance, the fulfillment of inequality (16) is uncertain. Larger parameter values have a lower likelihood of yielding improvements, although such improvements can be substantial, whereas smaller parameter values increase the likelihood of improvement, albeit potentially marginal or negligible, compared with the naïve forecast. Therefore, this boundary criterion highlights a critical trade-off between the likelihood of achieving improvements and the potential magnitude of these improvements.
In addition to using $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ as the estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, a more conservative approach can be used to generate the estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$:
  1. 1.
    Consider a data-generating process in which a sliding window, whose length is equal to that of the out-of-sample set, moves sequentially from the beginning to the end of the in-sample set. Each window can be treated as an "out-of-sample set" in the historical context.
  2. 2.
    Within each window, the directional accuracy is calculated and then translated into the corresponding value of $\theta$ via Eq. (9). As the window moves, a collection of $\theta$ values is generated, denoted as $\Theta_{\text{set}}$, whose bounds can be regarded as the empirical range of $\theta_{\text{out-of-sample}}^{*}$.
  3. 3.
    Similarly, a sliding window of the same length is applied to calculate a set of $\bar{\epsilon}$ values from the in-sample set, denoted as $\boldsymbol{\bar{\epsilon}}_{\text{set}}$, providing an empirical value range of $\bar{\epsilon}_{\text{out-of-sample}}$.
  4. 4.
    Using $\min(\Theta_{\text{set}})$ and $\min(\boldsymbol{\bar{\epsilon}}_{\text{set}})$ as conservative estimates for $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, the adjusted naïve forecast for time step $t$ in the out-of-sample set can be expressed as:
    $\hat{y}_{t}=y_{t-1}+\hat{D}_{t}\cdot\min(\Theta_{\text{set}})\cdot\min(\boldsymbol{\bar{\epsilon}}_{\text{set}})$ (17)
Algorithm 1 describes the process of generating the adjusted naïve forecast on the entire out-of-sample set.
Algorithm 1 Movement-Prediction-Adjusted Naïve Forecast on the Out-of-sample Set
Inputs:
$Y_{\text{in-sample}}$ //    Array of observations in the in-sample set
$Y_{\text{out-of-sample}}$ //    Array of observations in the out-of-sample set
$\hat{D}_{\text{in-sample}}$ //    Movement prediction on the in-sample set
$\hat{D}_{\text{out-of-sample}}$ //    Movement prediction on the out-of-sample set
$N$ //    Number of elements in the in-sample set
$M$ //    Number of elements in the out-of-sample set
Output:
$P$ //    Adjusted naïve forecast on the out-of-sample set
1. Initialize variables:
$D_{\text{in-sample}}\leftarrow[]$ //    Actual movements on the in-sample set
$P\leftarrow[]$
$ACC_{\text{set}}\leftarrow[]$ //    Directional accuracy set
$\Theta_{\text{set}}\leftarrow[]$ //    Theta set
$R_{\text{in-sample}}\leftarrow[]$ //    First-order difference set
$\boldsymbol{\bar{\epsilon}}_{\text{set}}\leftarrow[]$ //    Mean absolute increment set
2. Calculate actual movements:
for $t\leftarrow 1$ to $\text{length}(Y_{\text{in-sample}})$ do
     $\text{diff}\leftarrow Y_{\text{in-sample}}[t]-Y_{\text{in-sample}}[t-1]$ //    First-order difference
     $R_{\text{in-sample}}.\text{append}(\text{diff})$
     if $\text{diff}>0$ then
         $D_{\text{in-sample}}.\text{append}(1)$ //    Upward movement
     else
         $D_{\text{in-sample}}.\text{append}(-1)$ //    Downward or no movement
     end if
end for
3. Calculate directional accuracies:
for $t\leftarrow 1$ to $N-M+1$ do
@@ 3. Calculate directional accuracies:
     for $i\leftarrow t$ to $t+M-1$ do
         if $D_{\text{in-sample}}[i]==\hat{D}_{\text{in-sample}}[i]$ then
@@
         end if
     end for
     $ACC_{\text{set}}.\text{append}(\text{Correct\_Predictions}/M)$
end for
4. Compute the movement weights:
for each $\text{each }ACC\in ACC_{\text{set}}$ do
     $\theta\leftarrow 2\cdot\text{ACC}-1$
     $\Theta_{\text{set}}.\text{append}(\theta)$
end for
5. Compute the mean absolute increments:
for $t\leftarrow 1$ to $N-M+1$ do
     $\bar{\epsilon}\leftarrow\text{mean}(|R_{\text{in-sample}}[t:t+M-1]|)$
     $\boldsymbol{\bar{\epsilon}}_{\text{set}}.\text{append}(\bar{\epsilon})$
end for
6. Forecast the out-of-sample set:
$Y_{\text{last}}\leftarrow Y_{\text{in-sample}}[\text{end}]$ //    Start from the last element of the in-sample set
for $i\leftarrow 1$ to $\text{length}(Y_{\text{out-of-sample}})-1$ do
     $\Delta\leftarrow\min(\Theta_{\text{set}})\cdot\min(\boldsymbol{\bar{\epsilon}}_{\text{set}})\cdot\hat{D}_{\text{out-of-sample}}[i]$ //    Compute increment
     $Y_{\text{forecast}}\leftarrow Y_{\text{last}}+\Delta$ //    Compute the adjusted naïve forecast
     $P.\text{append}(Y_{\text{forecast}})$ //    Store result
     $Y_{\text{last}}\leftarrow Y_{\text{out-of-sample}}[i]$
end for
7. Return:
return $P$
3 Simulation
This section presents a simulation study using four synthetic time series to evaluate whether the adjusted naïve forecast outperforms the original naïve forecast when movement predictions with different directional accuracy levels are provided. The performance of the proposed method under each directional accuracy level was evaluated based on the out-of-sample set of each time series.
@@ 3.1 Datasets and Setup
The dataset includes four synthetic time series, each of which represents a symmetric random walk with a certain type of variance, as summarized in Table 1. Each time series, consisting of 2,500 steps, was generated using a fixed random seed (seed = 1) in the Python environment, as shown in Figure 2. To maintain all values within the positive range, an initial positive offset of 10000 units was applied to each series. The last 500 data points were selected as the out-of-sample set, and the remaining data points were used as the in-sample set. In addition, Figure 3 shows the first-order differences of each time series to visually demonstrate the properties of the increments, and Figure 4 shows the histograms of the first-order differences for all time series, highlighting the symmetric nature of the time series.
Table 1: Variance of synthetic random walks
Variance Type Expression Parameter Range
Constant variance $\sigma_{t}^{2}=\sigma_{0}^{2}$ $\sigma_{0}^{2}=1$
Linear variance trend $\sigma_{t}^{2}=\frac{\sigma_{0}^{2}}{1+k\cdot t}$ $\sigma_{0}^{2}=1$, $k\in(0,10)$
Cyclic variance $\sigma_{t}^{2}=\sigma_{0}^{2}(1+a\cdot\sin(\omega t))$ $\sigma_{0}^{2}=1$, $\omega=\frac{2\pi}{100}$, $a\in(0,10)$
Randomly varying variance $\sigma_{t}^{2}=\sigma_{t-1}^{2}+\eta_{t},\eta_{t}\sim\mathcal{N}(0,\xi^{2})$ $\sigma_{0}^{2}=1$, $\xi^{2}\in(0,1000)$
Figure 2: Synthetic time series: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
Figure 3: First-order differences of four synthetic time series: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
Figure 4: Histograms of first-order differences of four synthetic time series: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
@@
For each time series, the adjusted naïve forecast on the out-of-sample set was generated via Algorithm 1. The performance of each adjusted naïve forecast result was evaluated via four metrics: the RMSE, MAE, MAPE, and sMAPE. The RMSE, derived from the square root of the MSE, emphasizes larger errors, which makes it suitable for identifying significant deviations in forecasts. The MAE, on the other hand, provides an interpretable measure of the average error without overpenalizing outliers. The MAPE and sMAPE normalize the errors relative to the actual values, enabling comparisons between datasets with different scales. Together, these metrics capture both absolute and relative error characteristics, ensuring a robust and balanced evaluation. The formulas for these metrics are as follows:
@@
To statistically test the performance of the adjusted naïve forecast, for each synthetic time series at a specified directional accuracy level, the process to generate the adjusted naïve forecast was repeated 100 times. Each repetition uses a unique movement prediction sequence at the same directional accuracy level, which is generated by randomly replacing a subset of actual movements in the out-of-sample set with incorrect movements. The directional accuracy level varies in two ways: initially, it changes from 0.50 to 0.56 in increments of 0.01, and then it changes from 0.55 to 1.00 in larger increments of 0.05. This approach allows for finer granularity in the lower accuracy range, where small changes in the directional accuracy can have a significant effect on the performance of the method, while using larger steps at higher accuracy levels balances the computational efficiency. Consequently, for each synthetic time series, 100 independent values were calculated for each metric at each accuracy level.

### v9

@@ 2.2 Determination of the Optimal Coefficient $\theta$
Accordingly, the MSE difference between the naive forecast and the MPANF, denoted as $\Delta MSE_{\text{in-sample}}$, is expressed as follows:
$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{Naive}}-MSE_{\text{in-sample}}^{\text{MPANF}}$ (7)
@@
where $\theta_{\text{in-sample}}$ denotes the coefficient $\theta$ used by the MPANF on the in-sample set and $\bar{\epsilon}_{\text{in-sample}}$ denotes the mean of $|\epsilon_{t}|$ calculated over the in-sample set. A detailed algebraic derivation of the $\Delta MSE_{\text{in-sample}}$ is provided in Appendix A.
According to Eq. (7), maximizing $\Delta MSE_{\text{in-sample}}$ can further be reduced to maximizing the following function of $\theta_{\text{in-sample}}$:
@@
Since Eq. (8) is a concave function, differentiating it with respect to $\theta_{\text{in-sample}}$ and setting the derivative to zero yields the optimal solution:
Proposition 1 (Optimal coefficient). The optimal in-sample coefficient $\theta_{\text{in-sample}}^{*}$, which maximizes the in-sample MSE difference between the naive forecast and the MPANF, is given by:
@@
This equation indicates that the optimal coefficient $\theta$ depends solely on the in-sample movement prediction accuracy and is independent of the specific movement predictions or the characteristics of the time series.
@@
Proposition 2 (Performance improvement). Under the optimal coefficient setting using $\theta_{\text{in-sample}}^{*}$, the MPANF achieves a nonnegative improvement over the naive forecast in terms of the in-sample MSE:
@@
Proposition 2 confirms that as the directional accuracy increases beyond 0.5, the value of $\theta_{\text{in-sample}}^{*}$ increases from 0.0, leading to progressively greater improvements over the naive forecast. At a directional accuracy of 1.0, the maximum achievable improvement is $\bar{\epsilon}_{\text{in-sample}}^{2}$. Thus, with the optimal coefficient setting, the performance of the MPANF matches or surpasses that of the naive forecast on the in-sample set:
$MSE_{\text{in-sample}}^{\text{MPANF}}\leq MSE_{\text{in-sample}}^{\text{Naive}}$ (11)
2.3 Uncertainty Analysis of Out-of-Sample Forecasts
We now examine the out-of-sample performance of the MPANF. Suppose that the movement predictions on the out-of-sample set with $M$ time steps yield $m_{\text{out-of-sample}}^{\text{correct}}$ correct and $m_{\text{out-of-sample}}^{\text{incorrect}}$ incorrect predictions; then, the out-of-sample movement prediction accuracy, denoted as $ACC_{\text{out-of-sample}}$, can be expressed as:
@@ 2.3 Uncertainty Analysis of Out-of-Sample Forecasts
Suppose that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are used by the MPANF on the out-of-sample set; then, the out-of-sample MSE difference between the naive forecast and the MPANF, denoted as $\Delta MSE_{\text{out-of-sample}}$, is expressed as:
$\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{Naive}}-MSE_{\text{out-of-sample}}^{\text{MPANF}}$ (13)
@@
where $\bar{\epsilon}_{\text{out-of-sample}}$ denotes the out-of-sample mean absolute increment and $\theta_{\text{out-of-sample}}^{*}$ denotes the optimal out-of-sample coefficient $\theta$, which is derived according to Proposition 1:
@@
A detailed algebraic expansion of the $\Delta MSE_{\text{out-of-sample}}$ is provided in Appendix B. For the MPANF to outperform the naive forecast, the $\Delta MSE_{\text{out-of-sample}}$ should be no less than zero. Therefore, according to Eq. (13), the following inequality must hold:
@@
Given that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are the estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, respectively, inequality (15) can be generalized as:
Proposition 3 (Out-of-sample performance criterion). The necessary condition for the MPANF to achieve an improvement over the naive forecast on the out-of-sample set in terms of the MSE is as follows:
@@
where $\hat{\theta}_{\text{out-of-sample}}$ denotes the estimate of $\theta_{\text{out-of-sample}}^{*}$ and $\hat{\bar{\epsilon}}_{\text{out-of-sample}}$ denotes the estimate of $\bar{\epsilon}_{\text{out-of-sample}}$.
Proposition 3 establishes a boundary criterion when the MPANF can outperform the naive forecast. We can treat $\theta_{\text{out-of-sample}}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ as two random variables, each following a distribution defined over nonnegative values. Accordingly, inequality (16) represents a probabilistic event whose likelihood depends on the realized values of these two variables. Smaller values of these two variables generally increase the likelihood of improvement. However, as shown in Eq. (13), the resulting gains may be marginal. The MPANF can realize its maximum potential improvement only when $\hat{\theta}_{\text{out-of-sample}}$ and $\hat{\bar{\epsilon}}_{\text{out-of-sample}}$ are close to their respective true values, $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$.
In practice, the MPANF on the out-of-sample set can be derived via Algorithm 1.
Algorithm 1 Movement-prediction–adjusted naive forecast on the out-of-sample set
1:Inputs:
@@
[…]
6:Output:
7:$P$ $\triangleright$ Final point forecasts
8:Initialize:
[…]
10:Step 1: Determine $\theta$ and $\bar{\epsilon}$ from in-sample data
[…]
16:Step 2: Generate forecasts on the out-of-sample set
[…]
18:for $i\leftarrow 1,\dots,\text{length}(Y_{\text{out-of-sample}})$ do
19:  if $\hat{D}_{\text{out-of-sample}}[i]=1$ then
20:   $\Delta\leftarrow\theta\cdot\bar{\epsilon}$ $\triangleright$ Upward
21:  else
22:   $\Delta\leftarrow-\theta\cdot\bar{\epsilon}$ $\triangleright$ Downward
23:  end if
24:  $\hat{Y}_{i}\leftarrow Y_{\text{last}}+\Delta$
25:  $P.\text{append}(\hat{Y}_{i})$
26:  $Y_{\text{last}}\leftarrow Y_{\text{out-of-sample}}[i]$
27:end for
28:return $P$
3 Empirical Results
This section presents the empirical evaluation of the MPANF using real-world data. It includes a description of the datasets, details of the experimental setup, and a performance comparison of the proposed method against several baselines.
@@ 3.1 Datasets and Setup
In the empirical analysis, we first generate movement predictions for each target series and then apply the MPANF to produce the final point forecasts. Since developing a complex movement prediction method is beyond the scope of this study, the movement predictions are obtained using a simple co-movement–based approach with a selected exogenous variable, as detailed in Appendix C. This two-step design reflects the nature of the MPANF as a second-stage method that relies on accurate and consistent directional inputs.
Given the availability of data and the well-documented strong co-movement between the Financial Times Stock Exchange (FTSE) index and U.S. stock prices (Sarwar,, 2020; Gupta et al.,, 2024), together with the fact that the FTSE open price is recorded approximately 13 hours before the U.S. market close price on the same date, we selected eight U.S. financial time series as the target series, including Apple (AAPL), Boeing (BA), Google (GOOGL), Goldman Sachs (GS), Halliburton (HAL), Brent crude oil (OIL), Schlumberger (SLB), and Tesla (TSLA), based on their strong predictive alignment with FTSE movements. Accordingly, the FTSE open price is used as the exogenous variable for generating movement predictions.
Historical data for the FTSE open price and the close prices of the eight target series were obtained from Yahoo Finance, covering the period from October 8, 2014, to September 13, 2024. The FTSE data were aligned with each target series on a day-by-day basis, with extra observations removed and missing values imputed via forward filling (Junninen et al.,, 2004). After alignment, all series were standardized to a uniform length of 2,500 observations and then split into in-sample and out-of-sample sets at a 1:1 ratio, ensuring sufficient data for both model estimation and evaluation.The close prices of the eight target variables and the FTSE open price are illustrated in Figures 2 and 3, respectively, with a brief description provided in Table 1. The in-sample movement prediction accuracies for the eight target series all exceeded 0.55, indicating that the directional signals derived from the FTSE open price were meaningful.
Figure 2: Time series of eight target variables.
Figure 3: Open price of FTSE.
Table 1: Statistics of nine financial time series.
Time series Count Mean Std. Dev. Minimum Median Maximum $\mathbf{ACC}_{\text{in-sample}}$ $\mathbf{\bar{\epsilon}}_{\text{in-sample}}$
AAPL (Close price) 2500 88.46 62.00 20.72 54.16 234.55 0.5536 0.3842
BA (Close price) 2500 211.21 80.90 95.01 195.92 430.30 0.5568 2.5867
GOOGL (Close price) 2500 78.45 41.52 24.79 61.44 190.93 0.5504 0.4624
GS (Close price) 2500 239.35 89.70 117.58 202.27 510.25 0.5504 1.9030
HAL (Close price) 2500 30.99 9.61 4.32 33.30 151.09 0.5680 0.5252
OIL (Close price) 2500 66.43 18.59 19.33 65.20 127.98 0.5576 0.9067
SLB (Close price) 2500 45.94 15.45 11.02 49.49 74.91 0.5536 0.6596
TSLA (Close price) 2500 107.38 109.43 9.58 23.72 409.97 0.5536 0.3534
FTSE (Open price) 2500 7088.28 589.13 4993.90 7210.00 8445.80
Since many sophisticated models, such as machine learning or deep learning-based models, struggle to provide significant improvements over naive forecasts in financial time series forecasting, simple but strong baselines, such as naive, naive with drift, IMA (1,1), and linear regression, are selected for performance comparison Rogers, (1995); Jadevicius and Huston, (2015); Franses, (2020). In particular, the linear regression model, which acts as a meta-learner to combine the naive forecast and the movement prediction, uses only the current observed value and the predicted movement as predictors.
@@
The performance of each method was evaluated via four metrics: the RMSE, MAE, MAPE, and sMAPE. The RMSE, derived from the square root of the MSE, emphasizes larger errors, which makes it suitable for identifying significant deviations in forecasts. The MAE, on the other hand, provides an interpretable measure of the average error without over penalizing outliers. The MAPE and sMAPE normalize the errors relative to the actual values, enabling comparisons between datasets with different scales. Together, these metrics capture both absolute and relative error characteristics, ensuring a robust and balanced evaluation. The formulas for these metrics are as follows: